# To Buy an Electric Vehicle or Not? A Bayesian Analysis of Consumer Intent in the United States

### v1

@@ 5 Conclusion
EVs are increasingly regarded as a cornerstone of transportation decarbonization and broader climate change mitigation efforts. However, government support has varied between administrations. The Obama administration facilitated EV adoption by introducing federal tax credits for EV purchases and strengthening fuel economy and emissions standards, which incentivized automakers to develop electric models. In contrast, the Trump administration’s first term prioritized deregulation and rolled back these standards. The Biden administration implemented stricter emissions regulations and allocated substantial funding for EV infrastructure through the Bipartisan Infrastructure Law. A second Trump administration has signaled a likely return to deregulation, though specific policies remain unclear.
While policies and regulatory standards continue to evolve with changing administrations, the future of electric vehicles depends largely on the interest of consumers, which plays a crucial role in shaping national policies. Public opinion has a well-documented influence on US policymaking (Burstein, 2003; Shapiro, 2011), yet no prior research has systematically analyzed EV purchase sentiment using a nationally representative sample over multiple years or quantified the impact of the covariates. In addition, no studies have applied quantile analysis to examine how different factors influence the distribution of intention to purchase EVs.
This study addresses these gaps by examining public attitudes toward EVs using three nationally representative surveys conducted by the Pew Research Center from 2021-2023. Given the ordinal nature of the response data, we employ both ordinal probit regression and ordinal quantile regression to model EV purchase intent as a function of four novel variables not considered in existing studies, prior EV ownership, political affiliation, and demographic characteristics such as age, gender, income, and geographic location. We adopt the Bayesian approach to modeling and estimation, as it offers the entire posterior distribution of the parameters, enables direct probability statements, and allows us to account for uncertainty in observations and parameters when evaluating the effects of covariates.
The results from ordinal probit models indicate that respondents who say they have read or heard a lot about EVs, perceive them as environmentally beneficial, have previously owned an EV or hybrid vehicle, and are confident in the US developing charging infrastructure are more likely to express strong interest in purchasing an EV. Other factors positively influencing the intention to purchase an EV include being under 50 years old, having a higher level of education, living in metropolitan areas, residing in the West (compared to the Northeast), identifying with the Democratic Party (rather than the Republican Party), and having a higher income. In contrast, those who believe government climate efforts are excessive or adequate, are female, live in the Midwest, were born in the US, or are Black (as opposed to White) are less likely to express strong interest and more likely to report no interest in EVs. Results from covariate effects reveal that prior EV ownership, confidence in EV infrastructure, belief in EVs’ environmental benefits, attitudes toward government climate policy, and political affiliation have the strongest influence on individuals’ decision to purchase EVs.
@@
In addition, results from ordinal quantile analysis show that the relationship between these factors and EV purchase intent varies across the conditional distribution of the intention to purchase EVs. Respondents who are knowledgeable about EVs, believe in their environmental benefits, are confident in US charging infrastructure, have prior EV ownership, or identify as Democrats are especially likely to express strong purchase intent at higher quantiles of EV purchase intent. Similarly, those who believe EVs are environmentally beneficial, have confidence in infrastructure development, or identify as Democrats are particularly less likely to express no interest at lower quantiles. In contrast, respondents who believe the government is doing too much to address climate change are particularly more likely to express no interest at lower quantiles and particularly less likely to express strong interest at higher quantiles. Overall, the quantile models highlight significant variation in covariate effects across the distribution of EV purchase intent; thereby providing a more nuanced understanding than the ordinal probit model.
The findings demonstrate how targeted strategies may be able to help accelerate EV adoption. For consumers with low purchase intent (i.e., at lower quantiles), general awareness campaigns may have limited impact; instead, emphasizing practical aspects such as long-term cost savings, availability and reliability of charging infrastructure, and financial incentives could be more effective. For more interested consumers (i.e., at higher quantiles), reinforcing beliefs about environmental benefits can transform a likely customer to an actual EV owner. Confidence in charging infrastructure plays a dual role, both strengthening commitment among likely adopters and reducing outright rejection among skeptics, highlighting the potential value of visible investments and clear public communication about infrastructure progress. The results also indicate that political identity and views on climate policy are closely related to EV purchase intent, suggesting that framing EVs in terms of widely appealing features such as financial savings and energy independence—rather than solely as a climate solution—could resonate with a broader range of potential buyers. Moreover, for consumers wary of government climate policy, information provided by automakers, dealerships, and industry groups may carry more credibility than government-led campaigns. Overall, the empirical findings of our study provide useful insights for policymakers, industry stakeholders, and advertisers seeking to bolster adoption.

### v2

@@ 6 Conclusion
EVs are increasingly regarded as a cornerstone of transportation decarbonization and broader climate change mitigation efforts (Bhat et al., 2025). However, federal support has varied along partisan lines, with Democratic administration promoting EV adoption and Republican administration prioritizing deregulation and rolling back environmental standards. Amid these policy shifts, consumer interest remains a key driver of both market trends and national policy. Although public opinion is known to influence US policymaking (Burstein, 2003; Shapiro, 2011), yet no prior research has analyzed EV purchase sentiment using a nationally representative sample over multiple years or quantified the impact of covariates. In addition, no studies have applied quantile analysis to examine how different factors influence the distribution of intention to purchase EVs.
This study addresses these gaps by analyzing public attitudes toward EVs using three nationally representative Pew Research Center surveys from 2021-2023. We take the Bayesian route and model EV purchase intent with ordinal probit and ordinal quantile regression, incorporating novel predictors, prior EV ownership, political affiliation, and demographic factors. Results from the ordinal probit models show that strong interest in purchasing an EV is associated with greater exposure to EV information, belief in their environmental benefits, prior EV or hybrid ownership, and confidence in charging infrastructure. Additional positive predictors include being under 50, higher education and income, metropolitan residence, living in the West, and Democratic Party affiliation. Conversely, lower interest is linked to skepticism about government climate efforts, identifying as female or Black, living in the Midwest, and US-born status.
@@
Ordinal quantile analysis reveals that the effects of key factors on EV purchase intent vary across the distribution. Respondents who are knowledgeable about EVs, view them as environmentally beneficial, have confidence in charging infrastructure, own an EV, or identify as Democrats are more likely to express strong interest at higher quantiles of EV purchase intent. These same factors are associated with lower likelihood of no interest at lower quantiles. Conversely, those who believe the government is overreaching on climate policy are more likely to express no interest at lower quantiles and less likely to show strong interest at higher ones. These findings highlight how covariate effects differ across levels of purchase intent, thereby providing a more nuanced understanding than the ordinal probit model.
The findings offer valuable guidance for policymakers, industry stakeholders, and advertisers seeking to bolster EV adoption. Confidence in charging infrastructure is one of the strongest predictors of purchase intent—associated with greater interest among likely adopters and reduced rejection among skeptics—yet less than 20% respondents report high confidence. Infrastructure investment, along with clearer communication about availability, reliability and expansion plans, may help close this gap. Exposure to EV information is also linked to stronger intent and reduced disinterest, though only 37.7% of respondents report high exposure—highlighting the value of broader outreach through media, industry, and education. Belief that EVs are environmentally superior to gas-powered vehicles is similarly associated with higher intent and lower disinterest, with effects varying across the intent distribution; messaging focused on benefits such as improved air quality may therefore be effective. Perceptions that the government is doing too much on climate change are negatively associated with intent, suggesting that outreach emphasizing cost, convenience, or energy security—particularly from non-government sources—may have broader appeal.